OpenStudy (anonymous):

how do you solve: 2x+1= negative square root of 12

myininaya (myininaya):

since you raised both sides to an even power, you should check your results afterwards

myininaya (myininaya):

wait

myininaya (myininaya):

why did you square both sides/

OpenStudy (turingtest):

I don't know...:/

OpenStudy (turingtest):

lol

myininaya (myininaya):

subtract 1 on both sides first

myininaya (myininaya):

and then divide by 2 lol @ turing

myininaya (myininaya):

\[2x+1 =- \sqrt{12} => 2x=-1- \sqrt{12}=>x=\frac{-1-\sqrt{12}}{2}\]

myininaya (myininaya):

and there is a perfect square in 12 12=3(4)=3(2^2) so you can simplify more

OpenStudy (anonymous):

subtract 1, divide by 2

myininaya (myininaya):

yep lol